# A Unique Bosonic Symmetry in a 4D Field-Theoretic System

### v1

@@ 4 Bosonic Symmetry Transformations: Uniqueness Property and the Full Set of CF-Type Restrictions
In addition to our bosonic symmetry observations in (25) and (28), we note that the following bosonic transformations are true, namely;
$s_{\omega}\,{\cal L}_{(\bar{B})}=\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big\{\partial^{\mu}\,B^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,B^{\mu\nu}\big\}\,\bar{\cal B}_{\nu\sigma}$
$-\big\{\partial^{\mu}\,{\cal B}^{\nu\sigma}+\partial^{\nu}\,{\cal B}^{\sigma\mu}+\partial^{\sigma}\,{\cal B}^{\mu\nu}\big\}\,\bar{B}_{\nu\sigma}+B_{4}\,\partial^{\mu}B_{5}-B_{5}\,\partial^{\mu}B_{4}$
$-\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big)\,\bar{f}_{\nu}-\big(\partial^{\mu}\bar{F}^{\nu}-\partial^{\nu}\bar{F}^{\mu}\big)\,F_{\nu}\Big]+\partial_{\mu}\Big[\big(B_{1}\,\partial^{\mu}B-B\,\partial^{\mu}B_{1}\big)\Big]$
$+\;\partial_{\mu}\big[B_{\nu\sigma}+\bar{B}_{\nu\sigma}-(\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu})\big]\,\big(\partial^{\mu}\,{\cal B}^{\nu\sigma}+\partial^{\nu}\,{\cal B}^{\sigma\mu}+\partial^{\sigma}\,{\cal B}^{\mu\nu}\big)$
$-\;\partial_{\mu}\big[\bar{f}_{\nu}+\bar{F}_{\nu}-\partial_{\nu}\bar{C}_{1}\big]\,\big(\partial^{\mu}f^{\nu}-\partial^{\nu}f^{\mu}\big),$ (29)
$s_{\bar{\omega}}\,{\cal L}_{(B)}=\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big\{\partial^{\mu}\,\bar{\cal B}^{\nu\sigma}+\partial^{\nu}\,\bar{\cal B}^{\sigma\mu}+\partial^{\sigma}\,\bar{\cal B}^{\mu\nu}\big\}\,B_{\nu\sigma}$
$-\,\big\{\partial^{\mu}\,\bar{B}^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,\bar{B}^{\mu\nu}\big\}\,{\cal B}_{\nu\sigma}-B_{4}\,\partial^{\mu}B_{5}+B_{5}\,\partial^{\mu}B_{4}$
$+\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big)\,f_{\nu}+\big(\partial^{\mu}F^{\nu}-\partial^{\nu}F^{\mu}\big)\,\bar{F}_{\nu}\Big]+\partial_{\mu}\Big[\big(B\,\partial^{\mu}B_{1}-B_{1}\,\partial^{\mu}B\big)\Big],$
$-\;\partial_{\mu}\big[{\cal B}_{\nu\sigma}+\bar{\cal B}_{\nu\sigma}-(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu})\big]\,\big(\partial^{\mu}\,B^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,B^{\mu\nu}\big)$
$-\;\partial_{\mu}\big[f_{\nu}+F_{\nu}-\partial_{\nu}C_{1}\big]\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big),$ (30)
which establish that the coupled Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ respect both the bosonic symmetry transformations $s_{\omega}$ and $s_{\bar{\omega}}$ provided our whole theory is considered on the space of quantum fields where all the four CF-type restrictions are satisfied together. Furthermore, against the backdrop of our observations in (27), we note the following interesting observations for the transformations of the Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$, namely;
$\big(s_{\omega}+s_{\bar{\omega}}\big)\,{\cal L}_{(\bar{B})}=\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big\{\partial^{\mu}\,B^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,B^{\mu\nu}\big\}\,\bar{\cal B}_{\nu\sigma}$
$-\big\{\partial^{\mu}\,{\cal B}^{\nu\sigma}+\partial^{\nu}\,{\cal B}^{\sigma\mu}+\partial^{\sigma}\,{\cal B}^{\mu\nu}\big\}\,\bar{B}_{\nu\sigma}+B_{4}\,\partial^{\mu}B_{5}-B_{5}\,\partial^{\mu}B_{4}$
$-\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big)\,\bar{f}_{\nu}-\big(\partial^{\mu}\bar{F}^{\nu}-\partial^{\nu}\bar{F}^{\mu}\big)\,F_{\nu}\Big]+\partial_{\mu}\Big[\big(B_{1}\,\partial^{\mu}B-B\,\partial^{\mu}B_{1}\big)\Big]$
$+\;\partial_{\mu}\Big[B_{\nu\sigma}+\bar{B}_{\nu\sigma}-(\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu})\Big]\,\big(\partial^{\mu}\,{\cal B}^{\nu\sigma}+\partial^{\nu}\,{\cal B}^{\sigma\mu}+\partial^{\sigma}\,{\cal B}^{\mu\nu}\big)$
$-\;\partial_{\mu}\Big[\bar{f}_{\nu}+\bar{F}_{\nu}-\partial_{\nu}\bar{C}_{1}\Big]\,\big(\partial^{\mu}f^{\nu}-\partial^{\nu}f^{\mu}\big),$ (31)
$\big(s_{\omega}+s_{\bar{\omega}}\big)\,{\cal L}_{(B)}=\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big\{\partial^{\mu}\,\bar{\cal B}^{\nu\sigma}+\partial^{\nu}\,\bar{\cal B}^{\sigma\mu}+\partial^{\sigma}\,\bar{\cal B}^{\mu\nu}\big\}\,B_{\nu\sigma}$
$-\,\big\{\partial^{\mu}\,\bar{B}^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,\bar{B}^{\mu\nu}\big\}\,{\cal B}_{\nu\sigma}-B_{4}\,\partial^{\mu}B_{5}+B_{5}\,\partial^{\mu}B_{4}$
$+\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big)\,f_{\nu}+\big(\partial^{\mu}F^{\nu}-\partial^{\nu}F^{\mu}\big)\,\bar{F}_{\nu}\Big]+\partial_{\mu}\Big[\big(B\,\partial^{\mu}B_{1}-B_{1}\,\partial^{\mu}B\big)\Big],$
$-\;\partial_{\mu}\Big[{\cal B}_{\nu\sigma}+\bar{\cal B}_{\nu\sigma}-(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu})\Big]]\,\big(\partial^{\mu}\,B^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,B^{\mu\nu}\big)$
$-\;\partial_{\mu}\Big[f_{\nu}+F_{\nu}-\partial_{\nu}C_{1}\Big]]\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big).$ (32)

### v2

@@ 4 Bosonic Symmetry Transformations: Uniqueness Property and Full Set of CF-Type Restrictions
Hence, using the Gauss divergence theorem, it is clear that the action integral respects (i.e. $s_{\bar{\omega}}S=0$) the bosonic symmetry transformations ($s_{\omega}$) that have been listed in equation (26). In addition to our observations in equations (25) and (28), we note that the following bosonic symmetry transformation of the Lagrangian density ${\cal L}_{(\bar{B})}$ is also true, namely;
$s_{\omega}\,{\cal L}_{(\bar{B})}=+\,\partial_{\mu}\Big[\big(B_{1}\,\partial^{\mu}B-B\,\partial^{\mu}B_{1})\Big]+\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big\{\partial^{\mu}\,B^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,B^{\mu\nu}\big\}\,\bar{\cal B}_{\nu\sigma}$
$-\big\{\partial^{\mu}\,{\cal B}^{\nu\sigma}+\partial^{\nu}\,{\cal B}^{\sigma\mu}+\partial^{\sigma}\,{\cal B}^{\mu\nu}\big\}\,\bar{B}_{\nu\sigma}+B_{4}\,\partial^{\mu}B_{5}-B_{5}\,\partial^{\mu}B_{4}-\,\big(\partial^{\mu}f^{\nu}-\partial^{\nu}f^{\mu}\big)\,\bar{f}_{\nu}$
$-\big(\partial^{\mu}\bar{F}^{\nu}-\partial^{\nu}\bar{F}^{\mu}\big)\,F_{\nu}\Big]+\dfrac{1}{2}\Big[\big(\partial^{\mu}\bar{F}^{\nu}-\partial^{\nu}\bar{F}^{\mu}\big)\,\partial_{\mu}F_{\nu}+\big(\partial^{\mu}f^{\nu}-\partial^{\nu}f^{\mu}\big)\,\partial_{\mu}\bar{f}_{\nu}$
$+\big\{\partial^{\mu}\,{\cal B}^{\nu\sigma}+\partial^{\nu}\,{\cal B}^{\sigma\mu}+\partial^{\sigma}\,{\cal B}^{\mu\nu}\big\}\,\partial_{\mu}\bar{B}_{\nu\sigma}-\big\{\partial^{\mu}\,B^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,B^{\mu\nu}\big\}\,\partial_{\mu}\bar{\cal B}_{\nu\sigma}\Big],$ (29)
where the infinitesimal and continuous bosonic symmetry transformations ($s_{\omega}$) are listed in equation (24). On the other hand, the application of the bosonic symmetry transformation operator $s_{\bar{\omega}}$ [cf. Eq. (26)] on the Lagrangian density ${\cal L}_{(B)}$ leads to the following:
$s_{\bar{\omega}}\,{\cal L}_{(B)}=+\,\partial_{\mu}\Big[\big(B\,\partial^{\mu}B_{1}-B_{1}\,\partial^{\mu}B\big)\Big]+\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big\{\partial^{\mu}\,\bar{\cal B}^{\nu\sigma}+\partial^{\nu}\,\bar{\cal B}^{\sigma\mu}+\partial^{\sigma}\,\bar{\cal B}^{\mu\nu}\big\}\,B_{\nu\sigma}$
$-\,\big\{\partial^{\mu}\,\bar{B}^{\nu\sigma}+\partial^{\nu}\,\bar{B}^{\sigma\mu}+\partial^{\sigma}\,\bar{B}^{\mu\nu}\big\}\,{\cal B}_{\nu\sigma}+B_{5}\,\partial^{\mu}B_{4}-B_{4}\,\partial^{\mu}B_{5}+\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big)\,f_{\nu}$
$+\big(\partial^{\mu}F^{\nu}-\partial^{\nu}F^{\mu}\big)\,\bar{F}_{\nu}\Big]-\dfrac{1}{2}\Big[\big(\partial^{\mu}F^{\nu}-\partial^{\nu}F^{\mu}\big)\,\partial_{\mu}\bar{F}_{\nu}+\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big)\,\partial_{\mu}f_{\nu}$
$+\big\{\partial^{\mu}\,\bar{\cal B}^{\nu\sigma}+\partial^{\nu}\,\bar{\cal B}^{\sigma\mu}+\partial^{\sigma}\,\bar{\cal B}^{\mu\nu}\big\}\,\partial_{\mu}B_{\nu\sigma}-\big\{\partial^{\mu}\,\bar{B}^{\nu\sigma}+\partial^{\nu}\,\bar{B}^{\sigma\mu}+\partial^{\sigma}\,\bar{B}^{\mu\nu}\big\}\,\partial_{\mu}{\cal B}_{\nu\sigma}\Big].$ (30)
All the terms, that are outside the total spacetime derivatives, in the above equations (29) and (30), can be expressed in terms of all the four CF-type restrictions on our theory. To be precise, using a slightly involved algebraic exercise, we can recast the above equation (29), in terms of the CF-type restrictions, as follows:
$s_{\omega}\,{\cal L}_{(\bar{B})}=+\,\partial_{\mu}\Big[\big(B_{1}\,\partial^{\mu}B-B\,\partial^{\mu}B_{1})\Big]+\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big\{\partial^{\mu}\,B^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,B^{\mu\nu}\big\}\,\bar{\cal B}_{\nu\sigma}$ (31)
$-\big\{\partial^{\mu}\,{\cal B}^{\nu\sigma}+\partial^{\nu}\,{\cal B}^{\sigma\mu}+\partial^{\sigma}\,{\cal B}^{\mu\nu}\big\}\,\bar{B}_{\nu\sigma}+B_{4}\,\partial^{\mu}B_{5}-B_{5}\,\partial^{\mu}B_{4}-\,\big(\partial^{\mu}f^{\nu}-\partial^{\nu}f^{\mu}\big)\,\bar{f}_{\nu}$
$-\big(\partial^{\mu}\bar{F}^{\nu}-\partial^{\nu}\bar{F}^{\mu}\big)\,F_{\nu}\Big]+\frac{1}{2}\,\big[\partial^{\mu}\{f^{\nu}+F^{\nu}-\partial^{\nu}C_{1}\}-\partial^{\nu}\{f^{\mu}+F^{\mu}-\partial^{\mu}C_{1}\}\big]\,\partial_{\mu}\bar{f}_{\nu}$
$+\frac{1}{2}\,\big[\partial^{\mu}\{\bar{f}^{\nu}+\bar{F}^{\nu}-\partial^{\nu}\bar{C}_{1}\}-\partial^{\nu}\{\bar{f}^{\mu}+\bar{F}^{\mu}-\partial^{\mu}\bar{C}_{1}\}\big]\,\partial_{\mu}F_{\nu}-\frac{1}{2}\,\big[\partial^{\mu}\{B^{\nu\sigma}+\bar{B}^{\nu\sigma}$
$-(\partial^{\nu}\phi^{\sigma}-\partial^{\sigma}\phi^{\nu})\}+\,\partial^{\nu}\{B^{\sigma\mu}+\bar{B}^{\sigma\mu}-(\partial^{\sigma}\phi^{\mu}-\partial^{\mu}\phi^{\sigma})\}+\partial^{\sigma}\{B^{\mu\nu}+\bar{B}^{\mu\nu}$
$-(\partial^{\mu}\phi^{\nu}-\partial^{\nu}\phi^{\mu})\}\big]\,\partial_{\mu}\bar{\cal B}_{\nu\sigma}+\frac{1}{2}\,\big[\partial^{\mu}\{{\cal B}^{\nu\sigma}+\bar{\cal B}^{\nu\sigma}-(\partial^{\nu}\widetilde{\phi}^{\sigma}-\partial^{\sigma}\widetilde{\phi}^{\nu})\}+\,\partial^{\nu}\{{\cal B}^{\sigma\mu}+\bar{\cal B}^{\sigma\mu}$
$-(\partial^{\sigma}\widetilde{\phi}^{\mu}-\partial^{\mu}\widetilde{\phi}^{\sigma})\}+\partial^{\sigma}\{{\cal B}^{\mu\nu}+\bar{\cal B}^{\mu\nu}-(\partial^{\mu}\widetilde{\phi}^{\nu}-\partial^{\nu}\widetilde{\phi}^{\mu})\}\big]\,\partial_{\mu}\bar{B}_{\nu\sigma}.$
A close and careful look at the above equation demonstrates that, if we invoke the validity of all the four CF-type restrictions on our theory, we find that the r.h.s. of the above equation becomes a total spacetime derivative. In other words, the action integral (corresponding to the Lagrangian density ${\cal L}_{(\bar{B})}$) respects [cf. Eqs. (28),(31)] both the bosonic transformations $s_{\bar{\omega}}$ and $s_{\omega}$ on the subspace of quantum fields where all the four CF-type restrictions are satisfied. Against the backdrop of the above discussions, we observe that equation (30) can be recast, following exactly similar kind of algebraic exercise, in the following form:
$s_{\bar{\omega}}\,{\cal L}_{(B)}=+\,\partial_{\mu}\Big[\big(B\,\partial^{\mu}B_{1}-B_{1}\,\partial^{\mu}B\big)\Big]+\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big\{\partial^{\mu}\,\bar{\cal B}^{\nu\sigma}+\partial^{\nu}\,\bar{\cal B}^{\sigma\mu}+\partial^{\sigma}\,\bar{\cal B}^{\mu\nu}\big\}\,B_{\nu\sigma}$
$-\,\big\{\partial^{\mu}\,\bar{B}^{\nu\sigma}+\partial^{\nu}\,\bar{B}^{\sigma\mu}+\partial^{\sigma}\,\bar{B}^{\mu\nu}\big\}\,{\cal B}_{\nu\sigma}+B_{5}\,\partial^{\mu}B_{4}-B_{4}\,\partial^{\mu}B_{5}+\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big)\,f_{\nu}$
$+\big(\partial^{\mu}F^{\nu}-\partial^{\nu}F^{\mu}\big)\,\bar{F}_{\nu}\Big]-\frac{1}{2}\;\big(\partial^{\mu}\,B^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,B^{\mu\nu}\big)\,\partial_{\mu}\big[{\cal B}_{\nu\sigma}+\bar{\cal B}_{\nu\sigma}-(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu})\big]$
$+\frac{1}{2}\,\big(\partial^{\mu}\bar{F}^{\nu}-\partial^{\nu}\bar{F}^{\mu}\big)\,\partial_{\mu}\big[f_{\nu}+F_{\nu}-\partial_{\nu}C_{1}\big]+\frac{1}{2}\,\big[\partial^{\mu}\{B^{\nu\sigma}+\bar{B}^{\nu\sigma}-(\partial^{\nu}\phi^{\sigma}-\partial^{\sigma}\phi^{\nu})\}$
$+\,\partial^{\nu}\{B^{\sigma\mu}+\bar{B}^{\sigma\mu}-(\partial^{\sigma}\phi^{\mu}-\partial^{\mu}\phi^{\sigma})\}+\partial^{\sigma}\{B^{\mu\nu}+\bar{B}^{\mu\nu}-(\partial^{\mu}\phi^{\nu}-\partial^{\nu}\phi^{\mu})\}\big]\,\partial_{\mu}{\cal B}_{\nu\sigma}$
$-\frac{1}{2}\,\big[\partial^{\mu}\{\bar{f}^{\nu}+\bar{F}^{\nu}-\partial^{\nu}\bar{C}_{1}\}-\partial^{\nu}\big\{\bar{f}^{\mu}+\bar{F}^{\mu}-\partial^{\mu}\bar{C}_{1}\big\}\big]\,\partial_{\mu}f_{\nu}.$ (32)
We note, once again, that if we exploit the validity of the all the four CF-type restrictions on our theory, we observe that the r.h.s. of equation (32) becomes a total spacetime derivative. In other words, the action integral, corresponding to the Lagrangian density ${\cal L}_{(B)}$, respects [cf. Eqs. (25),(32)] both the bosonic symmetry transformations $s_{\omega}$ and $s_{\bar{\omega}}$ on the subspace of quantum fields where all the four CF-type restrictions are satisfied. Thus, in the language of the symmetry considerations, we have established that the coupled Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ respect both the bosonic symmetry transformations $s_{\omega}$ and $s_{\bar{\omega}}$ provided all the four CF-type restrictions are satisfied together.
We would like to add here that the operator equation: $(s_{\omega}+s_{\bar{\omega}})=0$ [cf. Eq. (27)], is satisfied at the level of the transformations of the coupled Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$, too, under the bosonic symmetry transformations. In this context, taking into account our observations in equations (28) and (31), it is very interesting to point out that the following is also true:
$\big(s_{\omega}+s_{\bar{\omega}}\big)\,{\cal L}_{(\bar{B})}=\dfrac{1}{2}\,\partial_{\mu}\,\Big[\Big(\partial^{\mu}\{B^{\nu\sigma}+\bar{B}^{\nu\sigma}-(\partial^{\nu}\phi^{\sigma}-\partial^{\sigma}\phi^{\nu})\}+\partial^{\nu}\{B^{\sigma\mu}+\bar{B}^{\sigma\mu}$
$-(\partial^{\sigma}\phi^{\mu}-\partial^{\mu}\phi^{\sigma})\}+\partial^{\sigma}\{B^{\mu\nu}+\bar{B}^{\mu\nu}-(\partial^{\mu}\phi^{\nu}-\partial^{\nu}\phi^{\mu})\}\Big)\,\bar{\cal B}_{\nu\sigma}-\Big(\partial^{\mu}\{{\cal B}^{\nu\sigma}+\bar{\cal B}^{\nu\sigma}$
$-(\partial^{\nu}\widetilde{\phi}^{\sigma}-\partial^{\sigma}\widetilde{\phi}^{\nu})\}+\partial^{\nu}\{{\cal B}^{\sigma\mu}+\bar{\cal B}^{\sigma\mu}-(\partial^{\sigma}\widetilde{\phi}^{\mu}-\partial^{\mu}\widetilde{\phi}^{\sigma})\}+\partial^{\sigma}\{{\cal B}^{\mu\nu}+\bar{\cal B}^{\mu\nu}$
$-(\partial^{\mu}\widetilde{\phi}^{\nu}-\partial^{\nu}\widetilde{\phi}^{\mu})\}\Big)\bar{B}_{\nu\sigma}-\,\big(\partial^{\mu}\{\bar{f}^{\nu}+\bar{F}^{\nu}-\partial^{\nu}\bar{C}_{1}\}-\partial^{\nu}\big\{\bar{f}^{\mu}+\bar{F}^{\mu}-\partial^{\mu}\bar{C}_{1}\big\}\big)\,F_{\nu}$
$-\,\big(\partial^{\mu}\{f^{\nu}+F^{\nu}-\partial^{\nu}C_{1}\}-\partial^{\nu}\{f^{\mu}+F^{\mu}-\partial^{\mu}C_{1}\}\big)\,\bar{f}_{\nu}\Big]+\dfrac{1}{2}\,\big(\partial^{\mu}\{f^{\nu}+F^{\nu}-\partial^{\nu}C_{1}\}$
$-\partial^{\nu}\big\{f^{\mu}+F^{\mu}-\partial^{\mu}C_{1}\big\}\big)\,\partial_{\mu}\bar{f}_{\nu}+\dfrac{1}{2}\,\big(\partial^{\mu}\{\bar{f}^{\nu}+\bar{F}^{\nu}-\partial^{\nu}\bar{C}_{1}\}-\partial^{\nu}\big\{\bar{f}^{\mu}+\bar{F}^{\mu}-\partial^{\mu}\bar{C}_{1}\big\}\big)\,\partial_{\mu}F_{\nu}$
$-\dfrac{1}{2}\,\Big[\Big(\partial^{\mu}\{B^{\nu\sigma}+\bar{B}^{\nu\sigma}-(\partial^{\nu}\phi^{\sigma}-\partial^{\sigma}\phi^{\nu})\}+\partial^{\nu}\{B^{\sigma\mu}+\bar{B}^{\sigma\mu}-(\partial^{\sigma}\phi^{\mu}-\partial^{\mu}\phi^{\sigma})\}$
$+\partial^{\sigma}\{B^{\mu\nu}+\bar{B}^{\mu\nu}-(\partial^{\mu}\phi^{\nu}-\partial^{\nu}\phi^{\mu})\}\Big)\Big]\,\partial_{\mu}\bar{\cal B}_{\nu\sigma}+\dfrac{1}{2}\,\Big[\Big(\partial^{\mu}\{{\cal B}^{\nu\sigma}+\bar{\cal B}^{\nu\sigma}-(\partial^{\nu}\widetilde{\phi}^{\sigma}-\partial^{\sigma}\widetilde{\phi}^{\nu})\}$
$+\partial^{\nu}\{{\cal B}^{\sigma\mu}+\bar{\cal B}^{\sigma\mu}-(\partial^{\sigma}\widetilde{\phi}^{\mu}-\partial^{\mu}\widetilde{\phi}^{\sigma})\}+\partial^{\sigma}\{{\cal B}^{\mu\nu}+\bar{\cal B}^{\mu\nu}-(\partial^{\mu}\widetilde{\phi}^{\nu}-\partial^{\nu}\widetilde{\phi}^{\mu})\}\Big)\Big]\,\partial_{\mu}\bar{B}_{\nu\sigma}.$ (33)
Thus, we observe that the operation of the sum (i.e. $s_{\omega}+s_{\bar{\omega}}$) of bosonic symmetry transportation operators $s_{\omega}$ and $s_{\bar{\omega}}$ on the perfectly anti-BRST invariant Lagrangian density ${\cal L}_{(\bar{B})}$ leads to the above transformation (33). A close look at all the individual terms on the r.h.s. of the above equation establish that all of them are expressed in terms of one of the total four CF-type restrictions on our theory. In other words, we have the validity of the operator relationship: $s_{\omega}+s_{\bar{\omega}}=0$ at the level of the coupled Lagrangian densities (as far as the symmetry considerations w.r.t. the bosonic symmetry transformation operators $s_{\omega}$ and $s_{\bar{\omega}}$ are concerned) provided we use all the four CF-type restrictions together. Against the backdrop of the above discussions, it is quite interesting to observe that the operation of the sum (i.e. $s_{\omega}+s_{\bar{\omega}}$) of the bosonic symmetry transformation operators on the perfectly BRST invariant Lagrangian density ${\cal L}_{(B)}$ is found to be as follows:
$\big(s_{\omega}+s_{\bar{\omega}}\big)\,{\cal L}_{(B)}=\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big(\partial^{\mu}\{\bar{f}^{\nu}+\bar{F}^{\nu}-\partial^{\nu}\bar{C}_{1}\}-\partial^{\nu}\big\{\bar{f}^{\mu}+\bar{F}^{\mu}-\partial^{\mu}\bar{C}_{1}\big\}\big)\,f_{\nu}$
$+\,\big(\partial^{\mu}\{f^{\nu}+F^{\nu}-\partial^{\nu}C_{1}\}-\partial^{\nu}\{f^{\mu}+F^{\mu}-\partial^{\mu}C_{1}\}\big)\,\bar{F}_{\nu}-\Big(\partial^{\mu}\{B^{\nu\sigma}+\bar{B}^{\nu\sigma}$
$-(\partial^{\nu}\phi^{\sigma}-\partial^{\sigma}\phi^{\nu})\}+\partial^{\nu}\{B^{\sigma\mu}+\bar{B}^{\sigma\mu}-(\partial^{\sigma}\phi^{\mu}-\partial^{\mu}\phi^{\sigma})\}+\partial^{\sigma}\{B^{\mu\nu}+\bar{B}^{\mu\nu}$
$-(\partial^{\mu}\phi^{\nu}-\partial^{\nu}\phi^{\mu})\}\Big)\,{\cal B}_{\nu\sigma}+\Big(\partial^{\mu}\{{\cal B}^{\nu\sigma}+\bar{\cal B}^{\nu\sigma}-(\partial^{\nu}\widetilde{\phi}^{\sigma}-\partial^{\sigma}\widetilde{\phi}^{\nu})\}+\partial^{\nu}\{{\cal B}^{\sigma\mu}+\bar{\cal B}^{\sigma\mu}$
$-(\partial^{\sigma}\widetilde{\phi}^{\mu}-\partial^{\mu}\widetilde{\phi}^{\sigma})\}+\partial^{\sigma}\{{\cal B}^{\mu\nu}+\bar{\cal B}^{\mu\nu}-(\partial^{\mu}\widetilde{\phi}^{\nu}-\partial^{\nu}\widetilde{\phi}^{\mu})\}\Big)\,B_{\nu\sigma}\Big]+\dfrac{1}{2}\;\big(\partial^{\mu}\bar{F}^{\nu}-\partial^{\nu}\bar{F}^{\mu}\big)\,$
$\;\partial_{\mu}\big[f_{\nu}+F_{\nu}-\partial_{\nu}C_{1}\big]-\dfrac{1}{2}\,\big(\partial^{\mu}\{\bar{f}^{\nu}+\bar{F}^{\nu}-\partial^{\nu}\bar{C}_{1}\}-\partial^{\nu}\big\{\bar{f}^{\mu}+\bar{F}^{\mu}-\partial^{\mu}\bar{C}_{1}\big\}\big)\partial_{\mu}f_{\nu}$
$+\dfrac{1}{2}\big(\partial^{\mu}\,B^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,B^{\mu\nu}\big)\partial_{\mu}\Big[{\cal B}_{\nu\sigma}+\bar{\cal B}_{\nu\sigma}-(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu})\Big]$
$+\dfrac{1}{2}\,\Big[\partial^{\mu}\{B^{\nu\sigma}+\bar{B}^{\nu\sigma}-(\partial^{\nu}\phi^{\sigma}-\partial^{\sigma}\phi^{\nu})\}+\partial^{\nu}\{B^{\sigma\mu}+\bar{B}^{\sigma\mu}-(\partial^{\sigma}\phi^{\mu}-\partial^{\mu}\phi^{\sigma})\}$
$+\,\partial^{\sigma}\{B^{\mu\nu}+\bar{B}^{\mu\nu}-(\partial^{\mu}\phi^{\nu}-\partial^{\nu}\phi^{\mu})\}\Big]\,\partial_{\mu}{\cal B}_{\nu\sigma}.$ (34)